1. BlackBox Co. had sale $10000 last year. The cost of goods sold was $6500, general and administrative expenses were $1000, interest expenses were $500, and depreciation was $1000. The firm’s tax rate is 35%.  Please calculate EBIT, Net Income and Cash flow from operation. [Hint: cash flow from operation is equal to net income plus depreciation expense because it is a non-cash expense and should be added back when calculating cash flow.]

2. During the last year of operation, account receivable increased by $10000, account payable decreased by $7000, and inventories decreased by $2000. What is the total impact of these changes on the difference between profit and cash flow?

3. Cisco Co. sells $1 billion marketable securities and use the money to buy inventories. What will happen to this firm’s current ratio?  How about its quick ratio?
Answer #1

1: a: EBIT = Sales – COGS- Selling and Administrative expenses-Depreciation

= 10000-6500-1000-1000

= $1500

B: Net Income = (EBIT- Interest)*(1-Tax)

= (1500-500)*(1-0.35)

=650

C: Cash flow from operation = Net Income+Depreciation

= 650+1000

=1650

2: There is no impact. The difference between profit and cash flow is due to depreciation which is not impacted by these variables.

3: Current ratio remains the same since current assets remain the same. Quick ratio will decline because quick assets are reduced.